# Accelerating Ill-conditioned Hankel Matrix RecoveryHanQin Cai, Longxiu Huang, Xiliang Lu, and Juntao You

### v1

@@ Abstract
keywords:
Ill-conditioned Hankel matrix, robust matrix completion, Newton-like descent
{MSCcodes}
15A29, 15A83, 47B35, 90C17, 90C26, 90C53
@@ 1 Introduction
$\mathcal{G}\bm{z}=\bm{X}\quad\textnormal{ and }\quad\mathcal{W}\bm{z}=[x_{1},x_{2},\cdots,x_{i},\cdots,x_{n-1},x_{n}]^{\top}.$
@@
(1) $\begin{split}\operatorname*{\mathrm{minimize}}_{{\bm{X}},{\bm{O}}\in\mathbb{C}^{n_{1}\times n_{2}}}&~{}\frac{1}{2p}\left\langle\mathcal{P}_{\bm{H}_{\Omega}}{\bm{Y}}-\mathcal{P}_{\bm{H}_{\Omega}}({\bm{X}}+{\bm{O}}),{\bm{Y}}-({\bm{X}}+{\bm{O}})\right\rangle\cr\operatorname*{\mathrm{subject~{}to}}&~{}{\bm{X}}\textnormal{ is rank-$r$ Hankel matrix and }{\bm{O}}\textnormal{ is sparse,}\end{split}$
@@
(2) $\begin{split}\operatorname*{\mathrm{minimize}}_{\bm{z},\bm{s}\in\mathbb{C}^{n}}\ &~{}\frac{1}{2p}\left\langle\Pi_{\Omega}\bm{f}-\Pi_{\Omega}(\bm{z}+\bm{s}),\bm{f}-(\bm{z}+\bm{s})\right\rangle\cr\operatorname*{\mathrm{subject~{}to}}&~{}\mathrm{rank}\left(\mathcal{G}\bm{z}\right)=r,\quad\Pi_{\Omega}\bm{s}\textnormal{ is }\alpha\textnormal{-sparse},\end{split}$
where $\Pi_{\Omega}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ is the observation operator on vector space that corresponds to the Hankel structured sampling pattern, $\bm{f}=\mathcal{G}^{*}{\bm{Y}}$ and $\bm{s}=\mathcal{G}^{*}{\bm{O}}$ are observed data and outlier matrices in their equivalent vector form, and $\alpha$ is a sparsity parameter defined later in Section 1.1. Notice that only the observed outliers will interfere with the recovery, thus we put the sparsity constraint on $\Pi_{\Omega}\bm{s}$ instead of the entire $\bm{s}$.
Since the applications of Hankel recovery are often time-sensitive, developing efficient solvers is a crucial task. Although many fast non-convex algorithms have been proposed in recent years, they often struggle with ill-conditioned problems (detailed in Section 1.3) which is one of the aforementioned major challenges in applications. This paper aims to study a provable non-convex algorithm for solving (2) that keeps high efficiency even if the condition number of the underlying Hankel matrix is large.
@@ 1.1 Assumptions
{assumption}
[Uniform sampling with replacement] On the vector space $\mathbb{C}^{n}$, the observation index set $\Omega=\{a_{k}\}_{k=1}^{m}$ is sampled uniformly with replacement from $[n]$. The observation rate is defined as $p:=|\Omega|/n=m/n$.
@@
{assumption}
[$\mu$-incoherence] Let ${\bm{U}}_{\star}\bm{\Sigma}_{\star}{\bm{V}}_{\star}^{*}$ be the compact singular value decomposition (SVD) of the underlying rank-$r$ Hankel matrix $\mathcal{G}\bm{z}_{\star}\in\mathbb{C}^{n_{1}\times n_{2}}$. Then $\mathcal{G}\bm{z}_{\star}$ is said $\mu$-incoherent if there exists a positive constant $\mu$ such that
@@
{assumption}
[$\alpha p$-sparsity] Given a sampling model (i.e., Section 1.1), the observed outliers are at most $\alpha p$-sparse. That is,
$\|\Pi_{\Omega}\bm{s}_{\star}\|_{0}\leq\alpha|\Omega|=\alpha pn.$
Sections 1.1, 1.1 and 1.1 have been widely used in matrix completion [14, 30, 35, 9], robust principal component analysis [13, 28, 3, 10], and low-rank Hankel recovery literature [18, 11, 12, 2]. Together, they ensure that (2) is a well-posed problem [15]. Note that Section 1.1 states only a $\alpha$ proportion of observed data are corrupted, without specifying the randomness of the corruption pattern. This condition holds with high probability as long as $\bm{s}_{\star}$ is $\frac{\alpha}{2}$-sparse [7]. Moreover, the sparsity of $\bm{s}_{\star}$ (resp. $\Pi_{\Omega}\bm{s}_{\star}$) implies the sparsity of $\mathcal{G}\bm{s}_{\star}$ (resp. $\mathcal{G}\Pi_{\Omega}\bm{s}_{\star}$) in each of its rows and columns, provided $n_{1}\approx n_{2}$.
@@ 1.3 Related work and contributions
The low-rank Hankel recovery problem has been extensively studied in the past decade. A naive approach is to directly apply some robust matrix recovery algorithms [19, 36, 20, 6, 7, 23, 5, 8] on the Hankel matrix which suffers from high computational complexities and inferior recoverability since the Hankel structure is ignored. Lately, many dedicated Hankel recovery algorithms have been proposed. Robust-EMaC [18] proposes a convex relaxation for the non-convex problem (1) and solves it with semidefinite programming which is still computationally too expensive. [11, 12, 33] propose three algorithms with linear convergence that solve the non-convex Hankel completion problem in $\mathcal{O}(r^{2}n+rn\log n)$ flops; however, none of them handles outliers and [33] also lacks theoretical guarantee. SAP [37] proposes to effectively solve (2) via alternating projections. The recovery guarantee with linear convergence is established with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{3}\log^{2}n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r))$. Note that iterative resample is required for SAP in theory. The computational complexity of SAP is $\mathcal{O}(r^{2}n\log n)$ flops and the accelerated version ASAP [2] improves it to $\mathcal{O}(r^{2}n+rn\log n)$ flops which however only studies the fully observed cases. HSGD [4] solves (2) via projected gradient descent in $\mathcal{O}(r^{2}n+rn\log n)$ flops. Although HSGD has a guaranteed linear convergence, it requires $\mathcal{O}(\kappa\log(1/\varepsilon))$ iterations to find a $\varepsilon$-solution, which is problematic for ill-conditioned problems. The theorem is established with $m\gtrsim\mathcal{O}(\max\{c_{s}^{2}\mu^{2}r^{2}\log n,c_{s}\mu\kappa^{3}r^{2}\log n\})$ and $\alpha\lesssim\mathcal{O}(1/(\max\{c_{s}\mu\kappa^{3/2}r^{3/2},c_{s}\mu r\kappa^{2}\}))$.
@@
    We propose a highly efficient algorithm for ill-conditioned robust Hankel recovery problems, coined Hankel Structured Newton-Like Descent (HSNLD). To find a $\varepsilon$-solution, its computational cost is $\mathcal{O}((r^{2}n+rn\log n)\log(1/\varepsilon))$ flops which beats all existing state-of-the-arts. Especially when compared with HSGD, our convergence rate is independent of the condition number.
@@
    The recovery guarantee, with linear convergence, has been established for HSNLD under some mild conditions. The sample complexity and outlier tolerance are $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$ respectively.
@@
    We verify the empirical advantages of HSNLD with synthetic data and real NMR signal. HSNLD achieves the best recoverability and computational efficiency in all ill-conditioned test problems. It gains more advantages when the condition numbers become even worse.
@@ 2 Proposed method
Hankel Structured Gradient Descent (HSGD) [4] is a state-of-the-art non-convex algorithm for Hankel matrix recovery problems. It has shown superior performance for well-conditioned Hankel matrices, both theoretically and empirically. However, when the underlying Hankel matrix is ill-conditioned, the convergence of HSGD is significantly slower in theory. In particular, HSGD requires $\mathcal{O}(\kappa\log\varepsilon^{-1})$ iterations to find a $\varepsilon$-solution. This phenomenon is empirically verified in Fig. 1, where we observe that HSGD suffers from the larger condition number of the underlying Hankel matrix. To overcome this major weakness of HSGD, we propose a novel non-convex algorithm inspired by recent developments on Newton-like Gradient Descent algorithms [33]. The proposed algorithm, dubbed Hankel Structured Newton-like Descent (HSNLD), is highly efficient even for ill-conditioned problems.
@@
(3) $\begin{split}\ell\left({\bm{L}},{\bm{R}};\bm{s}\right):=&~{}\frac{1}{2p}\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right),\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right\rangle\cr&~{}+\frac{1}{2}\left\lVert\left(\mathcal{I}-\mathcal{G}\mathcal{G}^{*}\right)\left({\bm{L}}{\bm{R}}^{*}\right)\right\lVert_{\mathrm{F}}^{2},\end{split}$
where $\bm{z}=\mathcal{G}^{*}{\bm{X}}\in\mathbb{C}^{n}$ and we parameterize the low-rank Hankel matrix ${\bm{X}}\in\mathbb{C}^{n_{1}\times n_{2}}$ as a product of two tall matrices ${\bm{L}}\in\mathbb{C}^{n_{1}\times r}$ and ${\bm{R}}\in\mathbb{C}^{n_{2}\times r}$, thus the low-rank constraint is implicitly encoded. Moreover, a regularization term $\left\lVert\left(\mathcal{I}-\mathcal{G}\mathcal{G}^{*}\right)\left({\bm{L}}{\bm{R}}^{*}\right)\right\lVert_{\mathrm{F}}^{2}$ is added to enforce the Hankel structure of ${\bm{L}}{\bm{R}}^{*}$. Targeting (3), the proposed HSNLD (summarized in Algorithm 1) will be discussed step by step in the following paragraphs. For the sake of presentation, the iterative steps will be discussed first, and then initialization.
Remark 2.1.
@@
1:Input: $\Pi_{\Omega}\bm{f}$: partial observation on the corrupted Hankel matrix in reweighted vector form; $r$: the rank of underlying Hankel matrix; $p$: observation rate; $\alpha$: outlier density; $\{\gamma_{k}\}$: parameters for sparsification operator; $K$: number of iterations;
@@
Iterative updates on $\bm{s}$. For the sake of computational efficiency, one should avoid forming the full-size Hankel matrix and process the calculation in the corresponding vector form. Hence, we first update $\bm{z}$ with the current low-rank components, i.e.,
(4) $\bm{z}_{k+1}=\mathcal{G}^{*}({\bm{L}}_{k}{\bm{R}}_{k}^{*}).$
By the definition of $\mathcal{G}^{*}$, the $t$-th entry of $\bm{z}$ can be written as
@@
Thus, (4) can be obtained via $r$ fast convolutions, with a low computational cost in $\mathcal{O}(rn\log n)$ flops.
To detect the outliers, we compute the residue $\Pi_{\Omega}(\bm{f}-\bm{z}_{k+1})$ over the observation and project it onto the space of sparse vectors via a sparsification operator. That is,
(5) $\bm{s}_{k+1}=\Gamma_{\gamma_{k}\alpha p}\Pi_{\Omega}\big{(}\bm{f}-\bm{z}_{k+1}\big{)},$
where $\Gamma_{\gamma_{k}\alpha p}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ is the vector sparsification operator that keeps the $\gamma_{k}\alpha pn$ largest-in-magnitude entries unchanged and vanishes the rest to 00. Therein, $\gamma_{k}>1$ is a parameter that tolerates slight overestimation on the outlier sparsity as the true $\alpha$ is usually unavailable to the user, and it also provides redundant robustness in outlier detection. We observe that taking $\gamma_{k}\rightarrow 1$ as $k\rightarrow\infty$ provides balanced performance in robustness and efficiency; however, $\gamma_{k}$ should always be larger than $1$. The computational cost of (5) is $\mathcal{O}(n\log n)$ flops or better.
Iterative updates on ${\bm{L}}$ and ${\bm{R}}$. The state-of-the-art HSGD [4] updates the low-rank Hankel components via iterative gradient descent. As discussed, the convergence of HSGD suffers from ill-conditioned problems. Inspired by the recent development in [33], we utilize an additional Newton-like preconditioner in the iterative updates of ${\bm{L}}$ and ${\bm{R}}$ to properly address the ill-conditioned Hankel matrices. We have
(6) $\begin{bmatrix}{\bm{L}}_{k+1}\\ {\bm{R}}_{k+1}\end{bmatrix}=\mathcal{P}_{\mathcal{C}}\begin{bmatrix}{\bm{L}}_{k}-\eta\nabla_{{\bm{L}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1}\\ {\bm{R}}_{k}-\eta\nabla_{{\bm{R}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{L}}^{*}_{k}{\bm{L}}_{k}\right)^{-1}\end{bmatrix}$
@@
for some numerical constant $C>0$. The choice of $C$ will be specified in Section 3 and the locally non-expansiveness of $\mathcal{P}_{\mathcal{C}}$ is verified in Lemma 5.12.
@@
$\quad~{}\nabla_{\bm{L}}\ell({\bm{L}},{\bm{R}};\bm{s})({\bm{R}}^{*}{\bm{R}})^{-1}$
$=\left[\mathcal{G}\big{(}p^{-1}\Pi_{\Omega}\left(\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right)\big{)}{\bm{R}}+\left({\bm{I}}-\mathcal{G}\mathcal{G}^{*}\right)({\bm{L}}{\bm{R}}^{*}){\bm{R}}\right]({\bm{R}}^{*}{\bm{R}})^{-1}$
$=\mathcal{G}\big{(}p^{-1}\Pi_{\Omega}\left(\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right)-\mathcal{G}^{*}({\bm{L}}{\bm{R}}^{*})\big{)}{\bm{R}}({\bm{R}}^{*}{\bm{R}})^{-1}+{\bm{L}}.$
One can compute $\bm{z}:=\mathcal{G}^{*}({\bm{L}}{\bm{R}}^{*})$ via $r$ fast convolutions which cost $\mathcal{O}(rn\log n)$ flops. Computing $p^{-1}\Pi_{\Omega}\left(\bm{z}+\bm{s}-\bm{f}\right)-\bm{z}$ costs $\mathcal{O}(n)$ flops or better. Since $\mathcal{G}(\cdot){\bm{R}}$ can be viewed as $r$ convolutions, it costs $\mathcal{O}(rn\log n)$ flops. Computing ${\bm{R}}({\bm{R}}^{*}{\bm{R}})^{-1}$ costs $\mathcal{O}(r^{2}n)$ flops, so does the projection $\mathcal{P}_{\mathcal{C}}$. Thus, the total computational cost of updating ${\bm{L}}$ (and similarly for ${\bm{R}}$) is $\mathcal{O}(rn\log n+r^{2}n)$
@@
(7) $\begin{split}\bm{s}_{0}&=\mathcal{W}^{-1}\Gamma_{\alpha p}\left(\mathcal{W}\Pi_{\Omega}\bm{f}\right),\cr{\bm{U}}_{0}\bm{\Sigma}_{0}{\bm{V}}_{0}^{*}&=\mathrm{SVD}_{r}(p^{-1}\mathcal{G}(\Pi_{\Omega}\bm{f}-\bm{s}_{0})),\cr\begin{bmatrix}{\bm{L}}_{0}\\ {\bm{R}}_{0}\end{bmatrix}&=\mathcal{P}_{\mathcal{C}}\begin{bmatrix}{\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2}\\ {\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2}\end{bmatrix}.\end{split}$
This initialization has shown promising performance and a low computational complexity of $\mathcal{O}(rn\log n)$. Note that the truncated SVD of a Hankel matrix has a low cost since all matrix-vector multiplications therein can be viewed as convolutions.
Therefore, the overall computational complexity of the proposed algorithm is $\mathcal{O}(rn\log n+r^{2}n)$ which is tied to the current state-of-the-art HSGD [4]. In addition, we emphasize that HSNLD never needs to form the entire Hankel matrix and that tracking the $n$ distinct elements in the corresponding vector form is sufficient. Thus, the proposed HSNLD is computationally and memory efficient.

### v2

# Accelerating Ill-conditioned Hankel Matrix Recovery via Structured Newton-like Descent

@@ Abstract
Keywords: Ill-conditioned Hankel matrix, robust matrix completion, Newton-like descent, precondition, direction of arrival, nuclear magnetic resonance.
@@ 1 Introduction
$\mathcal{G}\bm{z}=\bm{X}\quad\text{ and }\quad\mathcal{W}\bm{z}=[x_{1},x_{2},\cdots,x_{i},\cdots,x_{n-1},x_{n}]^{\top}.$ (1)
@@
$\begin{split}\operatorname*{\mathrm{minimize}}_{{\bm{X}},{\bm{O}}\in\mathbb{C}^{n_{1}\times n_{2}}}&~{}\frac{1}{2p}\left\langle\mathcal{P}_{\bm{H}_{\Omega}}{\bm{Y}}-\mathcal{P}_{\bm{H}_{\Omega}}({\bm{X}}+{\bm{O}}),{\bm{Y}}-({\bm{X}}+{\bm{O}})\right\rangle\cr\operatorname*{\mathrm{subject~{}to}}&~{}{\bm{X}}\textnormal{ is rank-$r$ Hankel matrix and }{\bm{O}}\textnormal{ is sparse,}\end{split}$ (2)
@@
$\begin{split}\operatorname*{\mathrm{minimize}}_{\bm{z},\bm{s}\in\mathbb{C}^{n}}\ &~{}\frac{1}{2p}\left\langle\Pi_{\Omega}\bm{f}-\Pi_{\Omega}(\bm{z}+\bm{s}),\bm{f}-(\bm{z}+\bm{s})\right\rangle\cr\operatorname*{\mathrm{subject~{}to}}&~{}\mathrm{rank}\left(\mathcal{G}\bm{z}\right)=r,\quad\Pi_{\Omega}\bm{s}\textnormal{ is }\alpha\textnormal{-sparse},\end{split}$ (3)
where $\Pi_{\Omega}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ is the observation operator on $\mathbb{C}^{n}$ that corresponds to the Hankel structured sampling pattern, $\bm{f}=\mathcal{G}^{*}{\bm{Y}}$ and $\bm{s}=\mathcal{G}^{*}{\bm{O}}$ are observed data and outlier matrices in their equivalent vector form, and $\alpha$ is a sparsity parameter defined later in Assumption 3. Since only the observed outliers interfere with the recovery, we put the sparsity constraint on $\Pi_{\Omega}\bm{s}$ instead of the entire $\bm{s}$.
Ill condition is one major challenge that remains unsolved for existing Hankel recovery literature, as the existing algorithms often struggle with convergence speed (detailed in Section 1.3). Unfortunately, ill-conditioned Hankel matrices often appear in many applications, specifically, when the frequency of a spectrally sparse signal is inseparable, the corresponding Hankel matrix has a large condition number [28, 29]. An example is Direction of Arrival (DOA) estimation which can be modeled as a low-rank Hankel recovery problem [39]. When multiple sources arrive from close angles, the corresponding Hankel matrix has a large condition number, see Figure 1 for illustration. DOA and many other applications of Hankel recovery, such as traffic anomaly detection [17] and seismic monitoring [16], are very time-sensitive, thus developing highly efficient solvers is an eager task. This paper aims to study a provable non-convex algorithm for solving (3) that keeps high computational efficiency even if the condition number of the underlying Hankel matrix is large.
Figure 1: Left: DOA estimation for $r=3$ far-field sources from the directions $\theta=87^{\circ},87.1^{\circ},87.3^{\circ}$, signal collected by a Uniform Linear Array with $n=2^{12}$ sensors. Right: Top six singular values of the corresponding square Hankel matrix. For a rank-3 approximation, the condition number $\kappa=\sigma_{1}/\sigma_{3}\approx 1083$. See Section 4.2 for more details of the DOA experiment.
@@ 1.1 Assumptions
Assumption 1 (Uniform sampling with replacement).
Let the observation index set $\Omega=\{a_{k}\}_{k=1}^{m}$ be sampled uniformly with replacement from $[n]$. The observation rate is $p:=|\Omega|/n=m/n$.
@@
Assumption 2 ($\mu$-incoherence).
Let ${\bm{U}}_{\star}\bm{\Sigma}_{\star}{\bm{V}}_{\star}^{*}$ be the compact singular value decomposition (SVD) of the underlying rank-$r$ Hankel matrix $\mathcal{G}\bm{z}_{\star}\in\mathbb{C}^{n_{1}\times n_{2}}$. Then $\mathcal{G}\bm{z}_{\star}$ is said $\mu$-incoherent if there exists a positive constant $\mu$ such that
@@
Assumption 3 ($\alpha p$-sparsity).
Given a sampling model (i.e., Assumption 1), the observed outliers are at most $\alpha p$-sparse. That is, $\|\Pi_{\Omega}\bm{s}_{\star}\|_{0}\leq\alpha|\Omega|=\alpha pn$.
Assumptions 1, 2 and 3 have been widely used in matrix completion [14, 33, 38, 9], robust principal component analysis [13, 31, 3, 10], and low-rank Hankel recovery literature [19, 11, 12, 2], ensuring that (3) is a well-posed problem [15]. Assumption 3 states that only a $\alpha$ proportion of observed data are corrupted, without specifying the randomness of the corruption pattern. This condition holds with high probability as long as $\bm{s}_{\star}$ is $\frac{\alpha}{2}$-sparse [7]. Moreover, the sparsity of $\bm{s}_{\star}$ (resp. $\Pi_{\Omega}\bm{s}_{\star}$) implies the sparsity of $\mathcal{G}\bm{s}_{\star}$ (resp. $\mathcal{G}\Pi_{\Omega}\bm{s}_{\star}$) in each of its rows and columns, provided $n_{1}\approx n_{2}$.
@@ 1.3 Related work and contributions
The low-rank Hankel recovery problem has been extensively studied in the past decade. A naive approach is to directly apply robust matrix recovery algorithms [18, 40, 20, 6, 7, 24, 5, 8] on the large $n_{1}\times n_{2}$ Hankel matrix which suffers from high computational complexities (i.e., $\mathcal{O}(n^{2}r)$ to $\mathcal{O}(n^{3})$ per iteration) and inferior recoverability since the Hankel structure is ignored. Lately, many dedicated Hankel recovery algorithms have been proposed. Robust-EMaC [19] proposes a convex relaxation for the non-convex problem (2) and solves it with semidefinite programming which is still computationally and memory too expensive. [11, 12, 36] propose three algorithms with linear convergence that solve the non-convex Hankel completion problem in $\mathcal{O}(r^{2}n+rn\log n)$ flops per iteration; however, none of them handles outliers and [36] also lacks theoretical guarantee for Hankel recovery problem. SAP [41] proposes to effectively solve (3) via alternating projections. The recovery guarantee is established with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{3}\log^{2}n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r))$. The overall computational complexity of SAP is $\mathcal{O}(r^{2}n\log n\log(1/\varepsilon))$ with a large hidden constant due to a truncated SVD used in every iteration. The accelerated version of SAP, namely ASAP [2], improves the complexity to $\mathcal{O}((r^{2}n+rn\log n)\log(1/\varepsilon))$ flops. However, ASAP studies only the fully observed cases. HSGD [4] solves (3) via projected gradient descent in $\mathcal{O}(r^{2}n+rn\log n)$ flops per iteration, with a small constant. Although HSGD has a guaranteed linear convergence, its iteration complexity is $\mathcal{O}(\kappa\log(1/\varepsilon))$, which is problematic for heavily ill-conditioned problems. The recovery guarantee of HSGD requires $m\gtrsim\mathcal{O}(\max\{c_{s}^{2}\mu^{2}r^{2}\log n,c_{s}\mu\kappa^{3}r^{2}\log n\})$ and $\alpha\lesssim\mathcal{O}(1/(\max\{c_{s}\mu\kappa^{3/2}r^{3/2},c_{s}\mu r\kappa^{2}\}))$.
@@
    We propose a highly efficient algorithm for heavily ill-conditioned robust Hankel recovery problems, coined Hankel Structured Newton-Like Descent (HSNLD). Its overall computational cost is $\mathcal{O}((r^{2}n+rn\log n)\log(1/\varepsilon))$ flops which beats all state-of-the-art algorithms. Especially when compared with HSGD, our convergence rate is independent of the condition number $\kappa$; when compared with SAP, our complexity is $\mathcal{O}(\min\{\log(n),r\})$ lower and the hidden constant is much smaller since no iterative truncated SVD involved.
@@
    The recovery guarantee, with linear convergence, has been established for HSNLD under some mild conditions. The sample complexity and outlier tolerance are $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$ respectively. Compared to HSGD, the balance regularization is omitted from the objective, a new distance measurement is employed and a series of new technical lemmas has been developed accordingly, which will benefit future studies in the field.
@@
    We verify the empirical advantages of HSNLD with synthetic datasets, as well as DOA and NMR signals. HSNLD achieves the best recoverability and computational efficiency in all ill-conditioned test problems. It gains more advantages when the condition numbers become even worse.
@@ 2 Proposed method
Hankel Structured Gradient Descent (HSGD) [4] is a state-of-the-art non-convex algorithm for Hankel matrix recovery problems. It has shown superior performance for well-conditioned Hankel matrices, both theoretically and empirically. However, when the underlying Hankel matrix is ill-conditioned, the convergence of HSGD is significantly slower in theory. In particular, HSGD requires $\mathcal{O}(\kappa\log\varepsilon^{-1})$ iterations to find a $\varepsilon$-solution. This phenomenon is empirically verified in Figure 2, where we observe that HSGD suffers from the larger condition number of the underlying Hankel matrix. To overcome this major weakness of HSGD, we propose a novel non-convex algorithm inspired by recent developments on Newton-like Gradient Descent algorithms [36]. The proposed algorithm, dubbed Hankel Structured Newton-like Descent (HSNLD), is highly efficient even for ill-conditioned problems.
@@
[…]
where $\bm{z}=\mathcal{G}^{*}{\bm{X}}\in\mathbb{C}^{n}$ and we parameterize ${\bm{X}}\in\mathbb{C}^{n_{1}\times n_{2}}$ as a product of ${\bm{L}}\in\mathbb{C}^{n_{1}\times r}$ and ${\bm{R}}\in\mathbb{C}^{n_{2}\times r}$, thus the low-rank constraint is implicitly encoded. Moreover, a regularization term $\left\lVert\left(\mathcal{I}-\mathcal{G}\mathcal{G}^{*}\right)\left({\bm{L}}{\bm{R}}^{*}\right)\right\lVert_{\mathrm{F}}^{2}$ is added to enforce the Hankel structure of ${\bm{L}}{\bm{R}}^{*}$. Targeting (4), the proposed HSNLD (summarized in Algorithm 1) will be discussed step by step in the following paragraphs. For the sake of presentation, the iterative steps will be discussed first, and then initialization.
Remark 1.
@@
1:Input: $\Pi_{\Omega}\bm{f}$: partial observation on the corrupted Hankel matrix in reweighted vector form; $r$: the rank of underlying Hankel matrix; $p$: observation rate; $\alpha$: outlier density; $\{\gamma_{k}\}$: parameters for sparsification operator; $\eta$: step size; $K$: maximum number of iterations.
@@
Iterative updates on $\bm{s}$. For the sake of computational efficiency, one should avoid forming the full-size Hankel matrix and process the calculation in the corresponding vector form. Hence, we first update $\bm{z}$ with the current low-rank components, i.e., $\bm{z}_{k+1}=\mathcal{G}^{*}({\bm{L}}_{k}{\bm{R}}_{k}^{*})$. By the definition of $\mathcal{G}^{*}$, the $t$-th entry of $\bm{z}$ can be written as
@@
Thus, $z_{k+1}$ can be obtained via $r$ fast convolutions, with a computational cost in $\mathcal{O}(rn\log n)$ flops.
To detect the outliers, we compute the residue $\Pi_{\Omega}(\bm{f}-\bm{z}_{k+1})$ over the observation and project it onto the space of sparse vectors via a sparsification operator. That is, $\bm{s}_{k+1}=\Gamma_{\gamma_{k}\alpha p}\Pi_{\Omega}\big{(}\bm{f}-\bm{z}_{k+1}\big{)}$, where $\Gamma_{\gamma_{k}\alpha p}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ is the vector sparsification operator that keeps the $\gamma_{k}\alpha pn$ largest-in-magnitude entries unchanged and vanishes the rest to 00. Therein, $\gamma_{k}>1$ is a parameter that tolerates slight overestimation on the outlier sparsity as the true $\alpha$ is usually unavailable to the user, and it also provides redundant robustness in outlier detection. We observe that taking $\gamma_{k}\rightarrow 1$ as $k\rightarrow\infty$ provides balanced performance in robustness and efficiency; however, $\gamma_{k}$ should always be larger than $1$. With off-the-shelf fast partial sorting algorithm, the computational cost of $s_{k+1}$ is $\mathcal{O}(n\log n)$ flops or better.
Iterative updates on ${\bm{L}}$ and ${\bm{R}}$. The state-of-the-art HSGD [4] updates the low-rank Hankel components via iterative gradient descent. As discussed, its convergence is hindered by ill-conditioning. Inspired by recent work [36, 10, 22], we utilize an additional Newton-like preconditioner in the iterative updates of ${\bm{L}}$ and ${\bm{R}}$ to properly address the ill-conditioned Hankel matrices. We have
$\begin{bmatrix}{\bm{L}}_{k+1}\\ {\bm{R}}_{k+1}\end{bmatrix}=\mathcal{P}_{\mathcal{C}}\begin{bmatrix}{\bm{L}}_{k}-\eta\nabla_{{\bm{L}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1}\\ {\bm{R}}_{k}-\eta\nabla_{{\bm{R}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{L}}^{*}_{k}{\bm{L}}_{k}\right)^{-1}\end{bmatrix}$ (5)
@@
for some numerical constant $C>0$. The choice of $C$ will be specified in Section 3 and the local non-expansiveness of $\mathcal{P}_{\mathcal{C}}$ is verified in Lemma 7.
@@
[…]
$=\mathcal{G}\big{(}\frac{1}{p}\Pi_{\Omega}\left(\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right)-\mathcal{G}^{*}({\bm{L}}{\bm{R}}^{*})\big{)}{\bm{R}}({\bm{R}}^{*}{\bm{R}})^{-1}+{\bm{L}}.$
One can compute $\bm{z}:=\mathcal{G}^{*}({\bm{L}}{\bm{R}}^{*})$ via $r$ fast convolutions which cost $\mathcal{O}(rn\log n)$ flops. Computing $\frac{1}{p}\Pi_{\Omega}\left(\bm{z}+\bm{s}-\bm{f}\right)-\bm{z}$ costs $\mathcal{O}(n)$ flops or better. Since $\mathcal{G}(\cdot){\bm{R}}$ can be viewed as $r$ convolutions, it costs $\mathcal{O}(rn\log n)$ flops. Computing ${\bm{R}}({\bm{R}}^{*}{\bm{R}})^{-1}$ costs $\mathcal{O}(r^{2}n)$ flops, so does the projection $\mathcal{P}_{\mathcal{C}}$. Thus, the total computational cost of updating ${\bm{L}}$ (and similarly for ${\bm{R}}$) is $\mathcal{O}(rn\log n+r^{2}n)$
@@
$\begin{split}\bm{s}_{0}=\mathcal{W}^{-1}\Gamma_{\alpha p}\left(\mathcal{W}\Pi_{\Omega}\bm{f}\right),{\bm{U}}_{0}\bm{\Sigma}_{0}{\bm{V}}_{0}^{*}=\mathrm{SVD}_{r}(\frac{1}{p}\mathcal{G}(\Pi_{\Omega}\bm{f}-\bm{s}_{0})),\begin{bmatrix}{\bm{L}}_{0}\\ {\bm{R}}_{0}\end{bmatrix}&=\mathcal{P}_{\mathcal{C}}\begin{bmatrix}{\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2}\\ {\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2}\end{bmatrix}.\end{split}$ (6)
This initialization has shown promising performance and a low computational complexity of $\mathcal{O}(rn\log n)$. Note that the truncated SVD of a Hankel matrix costs as low as $\mathcal{O}(rn\log n)$ flops since all matrix-vector multiplications therein can be viewed as convolutions.
Therefore, the overall computational complexity of the proposed algorithm is $\mathcal{O}(rn\log n+r^{2}n)$ which is tied to the current state-of-the-art HSGD [4]. In addition, HSNLD does not require forming the entire Hankel matrix, and only tracking the $n$ distinct elements in the corresponding vector form is sufficient. Thus, the proposed HSNLD is computationally and memory efficient.